Find the area of a cement walk 3 feet wide that surronds a rectangular plot of ground 86 feet long and 42 feet wide
Answer by checkley77(12844)

86*42=3,612 ft^2 area of the plot
(86+6)(42+6)=92*48=4,416 ft^2 area of the plot + the walk.
4,416-3,612=804 ft^2 is the area of the walk.